OpenStudy (anonymous):

help please Solve 2 log x = log 64 x = 8 x = ±8 x = 32 x = 128

OpenStudy (anonymous):

Raise both sides by then to get \(\LARGE 10^{2logx} = 64\). Remember exponentiation rules? 10^(ab) = (10^a)^b, so we get: \(\LARGE (10^{logx})^{2} = 64 -> x^2 = 64\) Can you conclude from this?

OpenStudy (anonymous):

ya i just wanted to double check my answer i got x=8 before thanks

OpenStudy (anonymous):

No problem :-)